OpenStudy (anonymous):

7x+5y=11 3x+y=7 solve for xy

OpenStudy (anonymous):

are you asking to solve for x*y or for the variables x and y

OpenStudy (anonymous):

to solve for the variables x and y, first you should choose to either solve this via substitution or elimination

OpenStudy (anonymous):

in this case, substitution will be easier. Original equations: 7x+5y=11 3x+y=7 first manipulate one so that you have solved for y: 3x-3x+y=7-3x y=-3x+7 then substitute in for y in the other equation (7x+5y=11) 7x+5(-3x+7)=11 7x-15x+35=11 -35 -35 -8x=-24 /-8 /-8 x=3 then plug your x value into one of the original equations and solve for y: 7(3)+5y=11 21+5y=11 5y=-10 y=-2

OpenStudy (anonymous):

of course you'll also want to double check answers by plugging them in

OpenStudy (anonymous):

7x+5y=11 3x+y=7 y = 7-3x I'm using substitution 7x + 5(7-3x)=11 7x + 35-15x = 11 -8x + 35 = 11 -8x = -24 x = 3 So... 7(3)+5y=11 21 + 5y = 11 5y = 11-21 5y = -10 y = -2 solution is (3,-2) no check solution in 2nd equation: 3x+y=7 3(3)+-2 = 7 9+-2=7 7=7 it works